What equal charges have to be placed on earth and moon to neutralise their gravitational attraction?

Grade:12

2 Answers

Vikas TU
6 years ago
For the Equilbrium ,
Fe = Fg
Kq^2/ r^2 = Gm1m2/r^2
q = charge
r = distance
m1 = mass of earth
m2  = mass of moon
q^2 = Gm1m2
on substituting the value
q = 0.86 x10^!4 C
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

For the Equilbrium ,
Fe = Fg Kq^2/ r^2
= Gm1m2/r^2
q = charge
r = distance
m1 = mass of earth
m2 = mass of moon
q^2 = Gm1m2 on substituting the value
q = 0.86 x 10^!4 C

Thanks and Regards